OpenStudy (anonymous):

Solve, showing all steps: 3 - 5x > -8

OpenStudy (anonymous):

I couuld guide you on how to solve it.

OpenStudy (anonymous):

I'm sorry. If I do that, I could get reported. Believe me, I would but rules are rules.

OpenStudy (anonymous):

\[3-5x >8\] -8 to both sides of the equation. 5/5=-5x/5 x = ?

OpenStudy (anonymous):

Solve that and you got your answer.

OpenStudy (anonymous):

Tell me what you got before you leave.

OpenStudy (anonymous):

-8-3=-11 -11/-5 x<2.2

OpenStudy (anonymous):

Is that your new equation or is that your answer?

OpenStudy (anonymous):

anwser

OpenStudy (anonymous):

The sign is correct. It is less than. However, the answer is incorrect. Solve this: \[\frac{ 5 }{ 5 }=\frac{ -5x }{ 5 }\]

OpenStudy (anonymous):

Basic cross multiplication. Tell me what you get.

OpenStudy (anonymous):

@rashional

OpenStudy (anonymous):

25=-25

OpenStudy (anonymous):

You know how to cross multiply correct? That's incorrect. This is the answer: \[x<1\]

OpenStudy (anonymous):

I'm sorry this is the answer: \[x <-1\]

OpenStudy (anonymous):

@rashional

OpenStudy (anonymous):

thats the final anwser and is that the full way of solvin it cross mult-

OpenStudy (anonymous):

On what you did?

OpenStudy (anonymous):

what you did

OpenStudy (anonymous):

Yes, I kinda showed you the steps. I subtracted -8 to both sides of the equation and got 5=-5x. After that, divide -5x to 5, and you'll end up with \[x <-1\]

OpenStudy (anonymous):

and thats it

OpenStudy (anonymous):

You understand why I flipped the sign?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@rashional

OpenStudy (anonymous):

so after you did 5/5=-5x/5 you got x < -1

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

yes i get it now thanks

OpenStudy (anonymous):

Great! Don't forget about the medal by clicking "Best Response" Have a good day.